# A General Class of Model-Free Dense Precision Matrix Estimators

### v1

@@ Abstract
We introduce prototype consistent model-free, dense precision matrix estimators that have broad application in economics. Using quadratic form concentration inequalities and novel algebraic characterizations of confounding dimension reductions, we are able to: (i) obtain non-asymptotic bounds for precision matrix estimation errors and also (ii) consistency in high dimensions; (iii) uncover the existence of an intrinsic signal-to-noise—underlying dimensions tradeoff; and (iv) avoid exact population sparsity assumptions. In addition to its desirable theoretical properties, a thorough empirical study of the S&P 500 index shows that a tuning parameter-free special case of our general estimator exhibits a doubly ascending Sharpe Ratio pattern, thereby establishing a link with the famous double descent phenomenon dominantly present in recent statistical and machine learning literature.
Keywords: Precision Matrix Estimation; Double Descent; Latent Space; Hidden Factors.
JEL Code: C18, C21, C55.
@@ 1 Introduction
In this paper, we propose a general class of new estimators for precision matrices in settings with a large—and potentially diverging—number of observed variables that is dense and model-free. Hence, we neither use penalized estimation nor estimate an imposed model to generate a precision matrix estimator.
Our main contributions are as follows. We introduce a general prototype class of consistent-non-sparse and unpenalized precision matrix estimators based on a deep connection between matrix algebra and hidden factor analysis. Hidden factor analysis is used only for achieving consistency in high dimensions via dimension reduction. Our estimators are model-free, and do not need to estimate hidden factors. Moreover, we derive non-asymptotic upper bounds for estimation errors that allow us to see the effects of a new complexity random variable, number of confounding factors, and signal-to-noise ratio. Our precision matrix estimator has three broad areas of interest. First, it can be used to construct debiased estimators for parameters in high-dimensional regression models. Debiased estimators are essential for constructing valid confidence intervals in high-dimensional settings (see Caner and Kock (2018), van de Geer et al. (2014)). A key challenge in this context is obtaining an estimate of the inverse of the Gram matrix, as the sample second moment matrix is typically singular when the number of variables exceeds the sample size. Our estimator provides a viable alternative for estimating this inverse, and can be readily adapted for use in such models. Second, in network analysis, under normally distributed random outcome variables, conditional dependence between two random variables is represented as a nonzero term in the precision matrix of all outcome variables. Our method avoids imposing “artificial” conditional independence assumptions—i.e., forcing zero entries in the precision matrix that correspond to the absence of links between variables. Instead, it preserves all connections in the network, allowing some to be minor and others to be dominant. A direct application of such setting is the study of tax interactions between US states in the study of De Paula et al. (2024). They impose sparsity of their social interaction matrix i.e., zero interactions between states (see equations (8) and (10) of their paper). Third, in financial econometrics, our estimator can be used where precision matrices of asset returns play a central role in tasks such as optimal portfolio allocation, portfolio variance estimation, and Sharpe Ratio calculation. A detailed literature review and discussion of these three broad applications is provided in Section 4.4.
1.1 Brief Overview of Estimators
We start by providing a formula for population precision matrix that can tie each variable to the other variables in a system. Then, to bypass the high-dimensionality of the problem, our conceptual innovation is to express each variable as a linear function of common but hidden factors, and solve for the corresponding precision matrix estimator. In this sense, hidden factors act as confounders. We make no identifying assumptions about hidden factors, nor estimate them.
As special subcases of our general estimator, we obtain a new adaptive principal components (adaptive PCR from now on) estimator and the ridgeless regression estimator (RRE from now on). We prove consistency of both estimators when the number of outcomes exceeds the sample size. Empirically, we find that the Sharpe Ratio of portfolios constructed using the ridge-less regression estimator exhibits a double ascent pattern in the S&P 500. This is significant, as it demonstrates that profitable portfolios can be formed without the need to estimate any tuning parameters.
We compare two subcase estimators: the ridge-less regression estimator (RRE) and adaptive principal component regression (PCR). In finite samples, the RRE has one fewer bias term than adaptive PCR, giving it a potential advantage. However, asymptotically the ranking may reverse, since the bias in adaptive PCR depends on the number of factors, while the bias in RRE depends on sample size. For instance, because the number of factors is typically smaller than the sample size, adaptive PCR may outperform RRE in estimating individual rows of the precision matrix.
We support these insights through simulations and empirical analysis. Specifically, we find that our ridge-less precision matrix estimator achieves a Sharpe Ratio of 0.2203 for a large portfolio of 357 S&P 500 stocks during the out-of-sample period from January 2015 to December 2023, after accounting for transaction costs. The closest competitor is the nodewise regression-based approach of Callot et al. (2021), which attains a Sharpe Ratio of 0.2082.
1.2 Dimension Reduction
Why do we use a hidden factor setup? First of all, we show that precision matrix estimation is related to linear regression and that can be related to hidden factors (see Online Appendix B). Each economic or financial variable can be related to latent factors and this can be thought as confounding variables as in causality related literature such as in Kaltenpoth and Vreeke (2023), Angrist and Pischke (2009), Imbens and Rubin (2015). Also the true variables of interest may be measured with errors, and latent factors can represent the ”true” variables rather than the reported ones, especially in emerging markets. Variable such as culture, sentiment, market can be extremely valuable in certain macro, financial data sets. These hidden factors in a linear model can be used to explain international stock returns such as in Singapore, Malaysia, USA as seen in Brooks and Del Negro (2004).
Another key aspect of using hidden factors is their effectiveness as a tool for dimension reduction in high-dimensional econometrics. This stems from the fact that the number of observed outcomes typically exceeds the number of underlying hidden factors. In our paper, we allow both the number of outcomes and the number of hidden factors to grow with the sample size. The role and significance of latent factors and the estimation of their key components, factors and loadings, have been largely extended and further linked to asset pricing in Lettau and Pelger (2020), causal inference and estimation of counterfactuals or treatment effects in Athey et al. (2019); Athey et al. (2021); Yan and Wainwright (2024); Kallus et al. (2018), recommenders systems in Agarwal et al. (2023), macroeconomic forecasting in Stock and Watson (2002a, b, 2016), matrix completion and missing panel data in Xiong and Pelger (2023); Cahan et al. (2023); Bai and Ng (2021); Duan et al. (2024); Bryzgalova et al. (2023). Chen et al. (2021) discuss how latent factors can model important features of network data such as reciprocity, degree heterogeneity, and homophily. They apply these methods to the estimation of gravity equations of trade between countries, showing how unobserved latent factors significantly influence observed trade volumes, even after controlling for traditional gravity variables like distance. Excellent theoretical and practical algorithmic results obtained in all these works usually serve as a strong indication that there is indeed a very strong rationale for utilization of latent factors. Somewhat paradoxically, while we will rely on similar factor modeling concepts and underlying intuitions that they bring along, we will be able to design new and novel precision matrix estimators without actually focusing on directly estimating factors themselves.
1.3 A Brief Discussion of Sparsity in Economics
Our proposed class of estimators is broadly applicable across a range of settings, including those where the underlying data do not exhibit sparsity. For instance, Giannone et al. (2021) analyze six widely used datasets in macroeconomics, microeconomics, and finance, and find that dense models consistently outperform sparse ones in terms of out-of-sample predictive accuracy. These applications include: (i) forecasting U.S. economic activity, (ii) identifying determinants of economic growth across countries, (iii) explaining the decline in crime rates across U.S. states, (iv) analyzing judicial rulings on eminent domain across U.S. circuits, (v) predicting the U.S. equity premium, and (vi) accounting for the cross-sectional variation in U.S. stock returns.
In asset pricing, the assumption of sparsity for explaining or predicting economic variables is particularly strong and has been critically examined (see p.33 and Section 3.7 of Nagel (2021)). In that study, sparsity is imposed between the outcome variable and the predictors, effectively forcing some coefficients to be zero—even if those predictors have a non-negligible but small relationship with the outcome. This can be problematic in settings where many variables contribute modestly to predictive performance. Techniques like Lasso may shrink all small coefficients to zero, resulting in an artificially sparse model and a significant loss of predictive power. [^1] Figure 3.2 of Nagel (2021) illustrates how sparsity-based methods can underperform due to this limitation.
1.4 Literature Survey
Several studies in the statistics literature have studied precision matrix estimation. An excellent survey is by Cai et al. (2016). In Section 3 of their paper, they provide two major approaches to precision matrix estimation. The first approach is neighborhood-based and closely related to nodewise regression, where a lasso regression is performed for each variable against all others. Alternatives to lasso, such as the Dantzig selector, can also be used within this framework (see Cai and Liu (2011), Callot et al. (2021)). The second major approach is based on penalized likelihood estimation. As shown by Bertsimas et al. (2020), lasso-type penalties in this context can induce shrinkage bias and lead to overly sparse precision matrix estimates, with many zero entries in each row. In a penalized likelihood based setting, they reformulate the problem as a mixed-integer optimization problem, which is solvable using convex optimization techniques. This approach produces better results than lasso based approaches. Sparsity of precision matrix is used in these two main approaches.
@@ 1.4 Literature Survey
A distinct approach, commonly used in finance and statistics, derives the precision matrix by inverting the covariance matrix using the Sherman–Morrison–Woodbury (SMW) formula. In this framework, sparsity is assumed in the covariance matrix of the idiosyncratic errors, while asset returns are modeled through a factor structure (see Fan et al. (2013), Caner et al. (2023)).
Our approach is different from the three approaches discussed above, and provides a new way of solving the same problem. Compared to the first two, we are not assuming exact sparsity in the precision matrix of outcomes. Compared to the last one, we are not using SMW formula to invert the covariance matrix. This allows us to bypass imposing a structure on the outcomes utilized in the covariance inversion. Instead we use a new precision matrix formula without sparsity and rely on hidden factors to generate each outcome. This effectively ensures that we do not benefit from the SMW matrix inversion formula. To be clear, we use model-free dense estimation, but consistency in high dimensions is achieved through hidden factor based dimension reduction without a need to estimate factors or identify them.
Note that nodewise regression based estimator of Callot et al. (2021) uses sparsity of precision matrix of outcomes. Then residual nodewise regression estimator of Caner et al. (2023) crucially depends on observed factors, and imposes sparse precision matrix of errors. Technique used in those papers cannot extend to hidden factors since observed factor structure in Caner et al. (2023) uses OLS estimation which is not applicable to more complicated estimation problems such as hidden factor structure. Note that Callot et al. (2021) has no factor structure and uses lasso based regression to estimate the precision matrix of outcomes.
We also tie our results to signal-to-noise ratio aspect which is not available in Callot et al. (2021), Caner et al. (2023) papers. We show that to get consistency of our estimators, signal-to-noise ratio should be larger than number of factors. Our method can be seen as a complementary approach to nodewise regression of Meinshausen and Bühlmann (2006). Nodewise regression uses an $l_{1}$ norm based penalization to form the rows of the precision matrix. Doing so, the underlying assumption is the imposition of sparsity in the precision matrix which may be interpreted as a strong assumption.
An excellent paper by Fan et al. (2013) provides a statistical high-dimensional analysis of covariance matrix of outcomes via hidden factors based estimation. Similar to us, it also relies on hidden factors but uses a thresholding based principal components estimator for the covariance matrix, and then inverts this through SMW formula. The key difference is that our estimator is model-free and uses hidden factors for consistency proof only. There are also other technical and conceptual differences that are worth highlighting. First, our proposed class of estimators is general and in addition to the PCR based ones it also includes the ridge-less regression estimator (RRE). Second, our signal assumption is much weaker as we do not need the minimum signal to grow at the same rate as the number of assets in the portfolio. Also, we allow for diverging noise assumption (this is a notable contrast to Fan et al. (2013) where the noise is tied to the maximum eigenvalue of the covariance matrix of errors with assumed constant dimension $p$), which is reflective of financial markets. Moreover, as the root of our approach resembles a finite sample analysis, it also allows to explicitly tie all the estimation errors to signal-to-noise ratio thereby making its importance in the rate of convergence estimates clearly visible. In particular, we uncover that the signal-to-noise ratio that dominates the number of factors ensures consistency of our precision matrix estimator.
@@
Our paper differs from Bing et al. (2021) and Bunea et al. (2022) in a fundamental way. These papers analyze general linear class of estimators in hidden factor models in linear regression model and is only interested in risk of the estimators. Their papers consider only one outcome variable and how it can be predicted with this new general class of estimators. Theorem 3 of Bing et al. (2021) specifically provides a risk upper bound and this bound depends on the magnitude of coefficients in a linear system. Our analysis on the other hand, considers the precision matrix estimation of outcome variable in linear regression via hidden factors (i.e. an application is large portfolio formation). We show how matrix algebra can be combined with latent factors to provide a new precision matrix formula in order to achieve consistency. Then a novel idea of estimation is to use a linear linkage between outcome variables to obtain precision matrix estimator based on the matrix algebraic new formula. One of the main difficulties, theoretically, is that the estimation of the main diagonal terms in the precision matrix estimation. The proof technique and ideas in Bing et al. (2021), Bunea et al. (2022) do not simply extend to this case. Hence, the precision matrix estimation is different than the risk analysis of a single outcome-asset. We also provide an explicit connection in our estimation error upper bounds to signal-to-noise ratio. The results in subcases are different also in a risk analysis versus precision matrix estimation. Bing et al. (2021), Bunea et al. (2022) show that in certain very high dimensional cases, ridge-less regression (RRE) can perform better than adaptive PCR in terms of risk asymptotically. However, in case of estimation errors of rows of precision matrix errors (in $l_{2}$ norm) we show adaptive PCR can perform better than RRE even in very high dimensional cases. These are due to different proofs for risk versus precision matrix estimation.
1.5 Notation
@@
We consider notation now. Define $C,c$ as generic positive constants. Note that transpose of any matrix $A:d\times l$ is represented by $A^{\prime}$, and transpose of a column vector, $a$ is represented by $a^{\prime}$ unless it is stated otherwise for notational convenience. Let $\|A\|_{2}$ represent the spectral norm of matrix $A$, and $\|a\|_{2}$ the Euclidean norm of a vector $a$. Also, we denote by $\|A\|_{F}$ the Frobenius norm of $A$. Let $\|A\|_{l_{\infty}}:=\max_{1\leq j\leq d}\|A_{j}\|_{1}$ be the maximum row sum norm, with $A_{j}:l\times 1$ as the $j$ th row among $d$ rows in matrix A, written in column format. The row format for $j$ th row of $A$ is $A_{j}^{\prime}:1\times l$. Let $\|A\|_{\infty}:=\max_{i,j}|A_{ij}|$, with $A_{ij}$ representing $i,j$ cell of matrix A. Denote the $k$ th largest singular value of $A$ matrix as $\sigma_{k}(A)$ with $k=1$ being the largest singular value. Let $A^{+}$ represent the Moore-Penrose inverse of matrix A. Let $\lambda_{k}(M)$ represent the $k$ th largest eigenvalue of $M:d\times d$, a square matrix, except $k=1,k=d$. However, for ease of understanding, the largest eigenvalue of $M$ is denoted as $Eigmax(M)$ and the minimum eigenvalue of $M$ is $Eigmin(M)$. $tr(M)$ represents the trace of matrix $M$. Let $a_{n}\asymp b_{n}$ show the same order for $a_{n},b_{n}$ random variables, i.e. $a_{n}=O_{p}(b_{n})$, and $b_{n}=O_{p}(a_{n})$. Section 2 proceeds with the idea of dimension reduction through hidden factors to achieve high dimensional consistency in the proofs. Section 3 introduces the estimators. Section 4 has consistency results. Section 5 provides empirical evidence. Conclusion shows future possible extensions of our work. Appendix shows the proof of Theorem 2. Simulations, some of the empirics and rest of the proofs are relegated to Online Appendix.
@@ 2 Dimension Reduction
In our setup we have $p$ outcome variables, and they are explained by $K$ common hidden factors to ensure dimension reduction, with $K<<p$. This will help in asymptotics and achieving consistency, however our estimators will be model-free and will not use hidden factors estimators. Our estimators will only use outcome variables and this will be clear in the next section. Specifically, our estimators are defined in (11)(12). Let $Y$ represent the $n\times p$ matrix of outcomes, each row representing the $t$ th time period, $t=1,\cdots,n$, and $j=1,\cdots,p$ representing the outcomes. $F$ represents $n\times K$ matrix of hidden factors, where each row, $t=1,\cdots,n$ represents a time period, and columns represent the hidden factors, $k=1,\cdots,K$. $A$ represents $K\times p$ matrix of factor loadings, and $U:n\times p$ as the error matrix. We have the following compact form of the linear hidden factor structure for dimension reduction (which will be decomposed via rows of Y into two equations immediately below in (2)-(3)).
$Y=FA+U.$ (1)
In accordance with our analysis, we subdivide all $p$ outcomes into $j$ th one and all the remaining ones. In that respect, start by defining each row of the outcome as $y_{t}:p\times 1$ (represents each row, in column form, whereas $y_{t}^{\prime}:1\times p$ is the row vector form) of $Y$ above, and $f_{t}:K\times 1$ represents each row (in column form) of $F$, and $u_{t}:p\times 1$ is the $t$ th row (in column form) of U. We decompose $A$ into $a_{j}:K\times 1$ factor loading for outcome j, and $A_{-j}:p-1\times K$ as factor loading matrix for all other outcomes except $j$. In other words $A:=[a_{j},A_{-j}^{\prime}]$.
We model $y_{jt}$ as the $j$ th outcome variable at time period $t=1,\cdots,n$. As an example, the $j$ th outcome here can be $j$ th asset’s excess return (over risk free rate such as 3 month T-Bill in USA) at time $t$. We define all the outcomes except $j$ th outcome as $Y_{-jt}:p-1\times 1$ vector at time $t$.
We decompose (1) into $j$ th outcome and the rest. Keeping all of the above in mind, an interesting parallel with the hidden (latent) factor regression in Bunea et al. (2022) and Bing et al. (2021) can be made as well.222 There in (1) of Bing et al. (2021) they use regressors to predict the outcome, and all of the regressors and outcome is a linear function of hidden variables. In the $t$ th row of $Y$ in (1), $j$ th outcome can be written as
@@
We use (2)(3) in our proofs rather than (1), and in this way our proofs are clear and tractable. Common hidden factors affect the $j$ th outcome in (2) as well as all remaining outcome variables in (3). Let $u_{jt},U_{-jt}$ represent the zero-mean errors for the $j$ th outcome and all other remaining outcomes, respectively.
In vector form, the j-th column of $Y$ in (1) can be written as
@@
where $Y_{-j}:n\times p-1$ matrix, and $U_{-j}:n\times p-1$ matrix too. As we will see later on, the vector form representation given in (4)(5) will turn out to be beneficial in our proofs as well. Our assumptions are stated next. Before doing so, we briefly introduce a few additional notational conventions.
@@
Define the variance of $j$ th error as $var(u_{jt})=\sigma_{j}^{2}\geq c>0$,where $c$ is a positive constant at time $t$, for $j=1\cdots,p$. Define the variance matrix of errors except the $j$ th one as $var(U_{-jt})=\Sigma_{U,-j}:p-1\times p-1$, for $j=1,\cdots,p$ at time $t$. So if $j=3$ for example $\sigma_{3}^{2}$ is the third asset error variance and $\Sigma_{U,-3}$ represents the covariance matrix of errors for all assets except the third one at time $t$. Define the covariance matrix of factors as $\Sigma_{f}:=cov(f_{t}):K\times K$. Define $y_{t}=(y_{jt},Y_{-jt}^{\prime})^{\prime}$ as the re-ordered $t$ th row of $Y$, and $\mu_{j}:=Ey_{jt}:1\times 1$, $\mu_{-j}:=EY_{-jt}:p-1\times 1$, and $\mu:=Ey_{t}:p\times 1$ at time $t$. Let the covariance matrix of outcomes be $\Sigma:=E(y_{t}-\mu)(y_{t}-\mu)^{\prime}$. Assumption 1 below will show iid nature of our data across $t=1,\cdots,n$. Define the precision matrix as $\Theta:=\Sigma^{-1}$. We provide Assumptions next.
Assumption 1.
(i). $y_{jt},Y_{-jt}$ are iid across $t=1,\cdots,n$ for each $j=1,\cdots,p$.
@@
(ii). $u_{jt},U_{-jt}$ are zero mean random variable and vector respectively, which are iid across $t=1,\cdots,n$, for each $j=1,\cdots,p$.
(iii). $f_{t}$ are iid across $t=1,\cdots,n$.
(iv). $f_{t},u_{jt},U_{-jt}$ are mutually independent across $t=1,\cdots,n$ given $j=1,\cdots,p$.
@@
Assumption 1 is a standard convention used in hidden factor literature (see, e.g., Bing et al. (2021) and Bunea et al. (2022).  [^3] Note also that (i)-(iii) implies that rows of F, Y, and U are iid. Note also, that independence of errors implies $Eu_{jt}U_{-jt}=0$ and $Eu_{t}u_{t}^{\prime}=I_{p},u_{t}=(u_{jt},U_{-jt}^{\prime})^{\prime}$ which is a sparsity restriction on the covariance matrix of errors. This is another assumption often present in the literature (see, e.g., Fan et al. (2008)) and amounts to a diagonal error covariance matrix. Relaxing this part of the assumption is certainly an interesting topic well worth of further exploring. However, since we pursue a general precision matrix of outcome variables in a hidden factor setup, and our outcomes are still correlated through the hidden factor structure, our extensive proofs would get more cumbersome. We therefore leave further work in this direction for a mathematically more narrowly focused future extension.
@@
Note that this sparsity for the error covariance matrix does not imply sparsity of the covariance matrix of outcome variables in our framework. We are not imposing sparsity on the links between outcome variables as done in nodewise regression context.
Assumption 2.

### v2

# Model-Estimation-Free, Dense, and High Dimensional Consistent Precision Matrix Estimators

@@ Abstract
Precision matrix estimation is a cornerstone concept in statistics, economics, and finance. Despite advances in recent years, estimation methods that are simultaneously (i) dense, (ii) consistent, and (iii) model-free are lacking. While each of these targets can be met separately, achieving them together is challenging.We address this gap by introducing a general class of estimators that unifies these features within a nonasymptotic framework, allowing for explicit characterization of the computational complexity–signal-to-noise ratio trade-off. Our analysis identifies three fundamental random quantities—complexity, signal magnitude, and method bias—that jointly determine estimation error. A particularly striking result is that ridgeless regression, a tuning-free special case within our class, exhibits the double descent phenomenon. This establishes the first formal precision matrix analogue to the well-known double descent behavior in linear regression. Our theoretical analysis is supported by a thorough empirical study of the S&P 500 index, where we observe a doubly ascending Sharpe ratio pattern, which complements the double descent phenomenon.
Keywords: Interpolation, Double Descent, Latent Space, Factor Models, Precision Matrix .
@@ 1 Introduction
Due to their usage in high dimensional portfolio formation in finance, network analysis in economics and debiased policy coefficient estimation in cross section-causality analysis, precision matrix estimation has been among the few cornerstone topics in statistics and finance in recent years.
We propose a general class of estimators that are simultaneously model estimation-free, dense, and consistent. Existing literature can handle separately one of these properties, but no precision matrix estimator can handle all of them jointly. Such a joint handling is highly desired as it is naturally expected to lead to better empirical performances.
Our design sharply differentiates our approach from classical shrinkage, factor-model, and nodewise-regression methods. For example, unlike shrinkage type of estimators, e.g., Ledoit and Wolf (2004), ours converge to precision matrix, not to a modified version of it. Compared to dense-model estimation free nonlinear shrinkage estimator of Ledoit and Wolf (2017), our estimator is high dimension consistent. Differently from the hidden factor model based estimator of , e.g., Fan et al. (2013), we do not have to impose identification, estimation, and rotation which allows for model free estimation. Finally, in sharp contrast to the nodewise regression, e.g., Callot et al. (2021), Meinshausen and Bühlmann (2006), we do not impose arbitrary sparsity assumption on the rows of precision matrix nor do we use penalized estimators that creates shrinkage bias. Our main contributions are summarized below.
(i). We introduce a general class of consistent-non-sparse and unpenalized precision matrix estimators. Linking a set of general dense-unpenalized estimators to precision matrix formula is new and novel. This is done without shrinkage, penalization or factor estimation. Hidden factor analysis is used only for achieving consistency via dimension reduction. (ii). We derive non-asymptotic upper bounds for estimation errors that allow us to see the effects of random variable that captures model complexity, number of confounding factors, and signal-to-noise ratio.(iii). Our class encompasses a broad range of precision matrix estimators, including ridgeless regression. For this estimator, we show the double descent due to the increasing size of the signal, when the number of outcomes exceeds the number of observations.(iv). Theoretical findings are complemented with very good empirical performance in S&P 500.
1.1 Literature Review
@@
Several studies in the statistics literature have studied precision matrix estimation. An excellent survey is by Cai et al. (2016) which shows neighborhood and penalization based approaches. We refer the readers to that survey for the literature and appendix that we have. A distinct approach, commonly used in finance and statistics, derives the precision matrix by inverting the covariance matrix using the Sherman–Morrison–Woodbury (SMW) formula. In this framework, sparsity is assumed in the covariance matrix of the idiosyncratic errors, while asset returns are modeled through a factor structure (see Fan et al. (2013), Caner et al. (2023)). Our approach is different from the approaches discussed in the literature. First of all, we are not assuming exact sparsity in the precision matrix of outcomes. We are also not using SMW formula to invert the covariance matrix.
Another approach is nodewise based. Note that nodewise regression based estimator of Callot et al. (2021) uses sparsity of precision matrix of outcomes. Then residual nodewise regression estimator of Caner et al. (2023) crucially depends on observed factors, and imposes sparse precision matrix of errors. Technique used in those papers cannot extend to hidden factors since observed factor structure in Caner et al. (2023) uses OLS estimation which is not applicable to more complicated estimation problems such as hidden factor structure. Recently, Caner and Daniele (2025) developed a new deep learning based precision matrix estimator. However, the estimator is consistent only when number of outcomes is less than or equal to sample size, which limits applications. We also tie our results to signal-to-noise ratio aspect which is not available in the studies of Callot et al. (2021), Caner et al. (2023). Our method does not impose sparsity like nodewise regression of Meinshausen and Bühlmann (2006). Nodewise regression uses an $l_{1}$ norm based penalization to form the rows of the precision matrix. Doing so, the underlying assumption is the imposition of sparsity in the precision matrix which may be interpreted as a strong assumption. An excellent paper by Fan et al. (2013) provides a statistical high-dimensional analysis of covariance matrix of outcomes via hidden factors based estimation. Similar to us, it also relies on hidden factors but uses a thresholding based principal components estimator for the covariance matrix, and then inverts this through SMW formula. The key difference is that our estimator is model estimation-free and uses hidden factors for consistency proof only.
@@
Our paper differs from Bing et al. (2021) and Bunea et al. (2022) in a fundamental way. These papers analyze general linear class of estimators in hidden factor models in linear regression model and is only interested in risk of the estimators. Their papers consider only one outcome variable and how it can be predicted with this new general class of estimators. Our analysis on the other hand, considers the precision matrix estimation of outcome variable in linear regression via hidden factors. In our paper, a novel idea is to use a linear linkage between outcome variables to obtain precision matrix estimator. One of the main difficulties, theoretically, is that the estimation of the main diagonal terms in the precision matrix estimation. The proof technique and ideas in Bing et al. (2021), Bunea et al. (2022) do not simply extend to this case.
Another set of estimators are shrinkage based ones. For example Ledoit and Wolf (2004) proposes a linear shrinkage estimator for covariances in iid data setting. This is to solve the invertibility problem of sample covariance matrix. Their estimator for covariance is dense, model-free hence making a major contribution. However, unlike our estimators, the estimator converges to a modified version of the covariance consistently not to the covariance matrix itself. Ledoit and Wolf (2017) propose a non-linear shrinkage estimator. This estimator specifically analyzes out-of-sample variance loss in iid data setting, and develop a modified covariance matrix estimator that increases the small eigenvalues, and decreases the large ones. It is a model-free dense estimator but the out-of-sample variance loss limit is different from the population one. Unlike our estimator, the precision matrix estimator in their paper is not consistent. But as mentioned, their main concern is different than consistency in high dimensions. More discussion about literature and hidden factors are in Appendix C.
@@
We consider notation now. Define $C,c$ as generic positive constants. Note that transpose of any matrix $Z:d\times l$ is represented by $Z^{\prime}$, and transpose of a column vector, $z$ is represented by $z^{\prime}$. Let $\|Z\|_{2},\|Z\|_{F}$ represent the spectral norm and Frobenius norm of matrix $Z$ respectively, and $\|z\|_{2}$ the Euclidean norm of a vector $z$. Let $\|Z\|_{l_{\infty}}:=\max_{1\leq j\leq d}\|Z_{j}\|_{1}$ be the maximum row sum norm, with $Z_{j}:l\times 1$ as the $j$ th row among $d$ rows in matrix Z, written in column format. The row format for $j$ th row of $Z$ is $Z_{j}^{\prime}:1\times l$. Let $\|Z\|_{\infty}:=\max_{i,j}|Z_{ij}|$, with $Z_{ij}$ representing $i,j$ cell of matrix Z. Denote the $k$ th largest singular value of $Z$ matrix as $\sigma_{k}(Z)$ with $k=1$ being the largest singular value. Let $Z^{+}$ represent the Moore-Penrose inverse of matrix Z. Let $\lambda_{k}(M)$ represent the $k$ th largest eigenvalue of $M:d\times d$, a square matrix, except $k=1,k=d$. However, for ease of understanding, the largest eigenvalue of $M$ is denoted as $Eigmax(M)$ and the minimum eigenvalue of $M$ is $Eigmin(M)$. $tr(M)$ represents the trace of matrix $M$. Section 2 proceeds with the idea of dimension reduction through hidden factors to achieve high dimensional consistency in the proofs. Section 3 introduces the estimators. Section 4 has consistency results. Section 5 provides empirical evidence. Conclusion shows future possible extensions of our work. Simulations, some of the empirics and all of the proofs, discussion about literature and hidden factors are relegated to Appendix.
@@ 2 Dimension Reduction
In our setup we have $p$ outcome variables, and they are explained by $K$ common hidden factors to ensure dimension reduction, with $K<<p$. This will help in asymptotics and achieving consistency, however our estimators will be model estimation-free and will not use hidden factors estimators. Our estimators will only use outcome variables and this will be clear in the next section. Specifically, our estimators are defined in (10)(11). Let $Y$ represent the $n\times p$ matrix of outcomes, each row representing the $t$ th time period, $t=1,\cdots,n$, and $j=1,\cdots,p$ representing the outcomes. $F$ represents $n\times K$ matrix of hidden factors, where each row, $t=1,\cdots,n$ represents a time period, and columns represent the hidden factors, $k=1,\cdots,K$. $A$ represents $p\times K$ matrix of factor loadings, and $U:n\times p$ as the error matrix. We have the following compact form of the linear hidden factor structure for dimension reduction (which will be decomposed via rows of Y into two equations immediately below in (2)-(3)).
$Y=FA^{\prime}+U.$ (1)
In accordance with our analysis, we subdivide all $p$ outcomes into $j$ th one and all the remaining ones. In that respect, start by defining each row of the outcome as $y_{t}:p\times 1$ (represents each row, in column form, whereas $y_{t}^{\prime}:1\times p$ is the row vector form) of $Y$ above, and $f_{t}:K\times 1$ represents each row of $F$, and $u_{t}:p\times 1$ is the $t$ th row of U. We decompose $A$ into $a_{j}:K\times 1$ factor loading vector for outcome j, and $A_{-j}:p-1\times K$ as factor loading matrix for all other outcomes except $j$. We model $y_{jt}$ as the $j$ th outcome variable at time period $t=1,\cdots,n$. We define all the outcomes except $j$ th outcome as $Y_{-jt}:p-1\times 1$ vector at time $t$. We decompose (1) into $j$ th outcome and the rest. In the $t$ th row of $Y$ in (1), $j$ th outcome can be written as
@@
We use (2)(3) in our proofs rather than (1), and in this way our proofs are clear and tractable. Common hidden factors affect the $j$ th outcome in (2) as well as all remaining outcome variables in (3). Let $u_{jt},U_{-jt}$ represent the zero-mean errors for the $j$ th outcome and all other remaining outcomes, respectively. In vector form, the j-th column of $Y$ in (1) can be written as
@@
where $Y_{-j}:n\times p-1$ matrix, and $U_{-j}:n\times p-1$ matrix too. As we will see later on, the vector form representation given in (4)(5) will turn out to be beneficial in our proofs as well.
@@
Define the variance of $j$ th error as $var(u_{jt})=\sigma_{j}^{2}\geq c>0$,where $c$ is a positive constant at time $t$, for $j=1\cdots,p$. Define the variance matrix of errors except the $j$ th one as $var(U_{-jt})=\Sigma_{U,-j}:p-1\times p-1$, for $j=1,\cdots,p$ at time $t$. Define the covariance matrix of factors as $\Sigma_{f}:=cov(f_{t}):K\times K$. Define $y_{t}=(y_{jt},Y_{-jt}^{\prime})^{\prime}$ as the re-ordered $t$ th row of $Y$, and $\mu_{j}:=Ey_{jt}:1\times 1$, $\mu_{-j}:=EY_{-jt}:p-1\times 1$, and $\mu:=Ey_{t}:p\times 1$ at time $t$. Let the covariance matrix of outcomes be $\Sigma:=E(y_{t}-\mu)(y_{t}-\mu)^{\prime}$. Assumption 1 below will show iid nature of our data across $t=1,\cdots,n$. Define the precision matrix as $\Theta:=\Sigma^{-1}$. We provide Assumptions next.
Assumption 1
@@
(i). $y_{jt},Y_{-jt}$ are iid across $t=1,\cdots,n$ for each $j=1,\cdots,p$. (ii). $u_{jt},U_{-jt}$ are zero mean random variable and vector respectively, which are iid across $t=1,\cdots,n$, for each $j=1,\cdots,p$. (iii). $f_{t}$ are iid across $t=1,\cdots,n$. (iv). $f_{t},u_{jt},U_{-jt}$ are mutually independent across $t=1,\cdots,n$ given $j=1,\cdots,p$.
@@
Assumption 1 is a standard convention used in hidden factor literature (see, e.g., Bing et al. (2021) and Bunea et al. (2022). [^1] Observe that (i)-(iii) implies that rows of F, Y, and U are iid. Note also, that independence of errors implies $Eu_{jt}U_{-jt}=0$ and $u_{t}=(u_{jt},U_{-jt}^{\prime})^{\prime}$ which is a sparsity restriction on the covariance matrix of errors. This is another assumption often present in the literature (see, e.g., Fan et al. (2008)) and amounts to a diagonal error covariance matrix. Relaxing this part of the assumption is certainly an interesting topic well worth of further exploring. Relaxing the iid structure of data for financial applications is a good idea, however at this point our proofs, via Assumption 4 below as well, use subgausssian-iid based quadratic form concentration inequalities in high dimensional probability literature. Having a different structure in the proofs and solving time series nature is a key issue that we can follow up in future directions tied to financial applications.
@@
We are not imposing sparsity on the links between outcome variables as done in nodewise regression context.
Assumption 2